The following table shows a partial listing of T-bills from the Wall Street Journal

Outline. Rather than providing prices of each bill, the financial press reports yields using the bank-discount method. The face value of T-bill is $10,000.


Treasury Bills

August 6, 2021

 Maturity  Bid (%)  Asked (%)

9/2/2021  0.040  0.030

9/9/2021  0.045  0.035

9/16/2021  0.043  0.033

9/23/2021  0.048  0.038


Consider the bill maturing on September 23.

a. What is the price at which the dealer is willing to buy assuming days to

maturity are 48 days?


b. What is the bond equivalent yield if you purchase the bill?


1
Expert's answer
2022-02-23T12:44:32-0500

a) Calculation of price to be charged by dealer for selling t-bill maturing on 9/23/2021:


The vender sells a t-bill at bid cost or rate to the purchaser (along these lines, normally bid is higher than ask, connoting more grounded dealer over its purchasers), while ask cost or rate, lower cost acknowledged by purchaser as against its sell.


Consequently, merchant will set the cost of t-charge utilizing bid rate, which is 0.04% of t-bill developing on 9/2/2021.


Bid (0.04%) is as indicated by 360 days, along these lines, first let us work out day by day rate by separating the bid rate by 360 days (changing over 0.09% in numbers, ("\frac{0.04} {100} = 0.0004" )


Every day pace of t-bill="\frac{0.0004} {360} = 0.0000011"


The t-charges are of 48 days. thusly 48 days returns would be 48 days * 0.0000011 = 0.0000528


Estimation of cost of t-charges: Face esteem * (1-returns of t-bills till development)


Price= "\$10,000 \times (1-0.0000528)"


In this manner, Price of t-bills developing on 9/2/2021= $9,999.472



b) Computation of Bond Equivalent Yield:


It empowers financial backers to register their annualized yield from the speculation and contrast something similar and those of customary securities.


Formula and calculation of Bond Equivalent yield is as per the following:


Bond Equivalent Yield = (Face Value - Purchase Price)/Purchase Price * 365days/n


="(\frac{10000-9999.472} {9999.472}) \times\frac{365} {48} =0.00040152"